The length of a rectangular garden is 10 feet longer than its width. If the garden's perimeter is 208 feet, what is the area of the garden in square feet?
Answer by Boreal(15235)

x and 104-x are the length and width; length=width= half the perimeter or 104 feet.
but x-10 = 104-x
2x=114
x=57 feet length
x-10=47 foot width
area is 57*47=2679 ft^2.